# Memory-Centric Embodied Question Answer

### v1

@@ 5.3 Ablation Studies
As shown in Table 2, the ablation study evaluates the impact of the injection of memory into various module on MT-HM3D and HM-EQA datasets. Without memory injection (None), baseline success rates are 34.3% (MT-HM3D) and 44.6% (HM-EQA), with suboptimal step efficiency. Introducing the Stop Criterion (S) alone yields marginal improvements in MT-HM3D (+0.8% Succ., reduced steps). Combining Stopping and Answering Module (S+A) significantly boosts performance: MT-HM3D’s success rises by 12.5% (to 46.8%), and HM-EQA improves by 7.6% (to 52.2%), confirming that memory-enhanced answering critically resolves ambiguities in both tasks. Full integration (S+A+P) achieves peak results: MT-HM3D reaches 51.2% Succ. (16.9% gain over baseline) with a 0.43 norm step reduction, while HM-EQA attains 59.9% success rate (+15.3%) and 25% fewer steps (0.52). The planner’s contribution is crucial, in synergy with S and A to optimize task decomposition and trajectory planning. This demonstrates that memory injection benefits all modules cumulatively, with the planner driving efficiency by mitigating redundant exploration. Notably, HM-EQA’s higher baseline and performance with S+A+P suggest its questions demand multi-phase reasoning, where memory-centric framework excels. In summary, the injection of memory into various modules is essential to improving success rates and efficiency in EQA tasks.
We have evaluated the impact of both global and local memory on the model’s performance. As shown in Table 3, both local and global memory play significant roles in problem-solving. Specifically, local memory enhances the success rate by 4.9%, while global memory boosts exploration efficiency by approximately 20%. From these findings, we can conclude that global memory retains the location and category of entities within the scene, aiding the planner in navigating directly to the target location and enabling the answering module to respond to questions based on past explorations. Local memory, on the other hand, preserves fine-grained information during the current exploration, allowing the model to answer questions with greater accuracy.
6 Limitations
In our implementation, the exploration process lacks interpretability and a basis for decision-making. Additionally, the model may generate responses without relying on the information obtained during exploration, and current evaluation metrics may inaccurately assess performance in such cases. In the future, we will build trajectory annotations based on the MT-HM3D, including thought processes, tool-chains, and the outputs of executing these tool-chains. Based on data with trajectory annotations, we will design an evaluation metric that takes the exploration process into account.
@@ 7 Conclusion
In this paper, we propose a memory-centric EQA framework, which contrasts with planner-centric frameworks. In our method, hierarchical memory information is feed into the planner, stop module, and answering module. The memory guides reasoning for various module, prevented redundant exploration and improved success rate of response in multi-target, multi-scenario tasks. Additionally, we have constructed a multi-target EQA dataset MT-HM3D based on the real-world HM3D scenes. This dataset includes four categories of multi-target tasks, allowing for comprehensive validation of the memory capabilities of EQA models. The effectiveness of MemoryEQA has been demonstrated through performance on the HM-EQA and MT-HM3D datasets.
@@
Appendix A Method Details
We utilize prompt 1 and prompt 2 to prompt VLM to generate coarse-grained scene description and fine-grained object description respectively.
@@ Appendix A Method Details
Prompt 1: Prompt for coarse-grained scene description.
@@
Prompt 2: Prompt for fine-grained object description.
@@ E. Very high
Prompt 3: Prompt for the evaluation of confidence.
Prompt 4 and prompt 5 are used to prompt the answering module to closed-vocabulary response and open-vocabulary response.
{Question} Answer with the option’s letter from the given choices directly.’
Prompt 4: Prompt for the response.
@@
Prompt 5: Prompt for the response on open-vocabulary data.
Next, prompt 6 and prompt 7 follow ExploreEQA [27] setting to prompt VLM to output explore direction.
@@
Prompt 6: Prompt for the explore direction.
@@
Prompt 7: Prompt for the Exploring Value.
Appendix B Data Construction
Figure 5: Statistic of our proposed MT-HM3D.
B.1 Statistics
To evaluate the memory capabilities of EQA models, we propose an EQA dataset that features multiple targets distributed across various regions. As shown in Table  4, our dataset exhibits the highest average number of targets and regions, significantly increasing the complexity of the task. The requirement to explore multiple targets across different regions poses a substantial challenge to the model’s memory capacity. The statistical details of the MT-HM3D dataset are illustrated in Figure  5.
Dataset Simulator Source Real Scenes Objects Region Creation
EQA-v1[11] House3D SUNCG - - Rule-Based
MP3D-EQA[34] MINOS MP3D - - Rule-Based
MT-EQA[39] House3D SUNCG - - Rule-Based
IQA[14] AI2THOR - - - Rule-Based
EXPRESS-Bench[16] Habitat HM3D - - VLMs
HM-EQA[27] Habitat HM3D 2.87 0.51 VLMs
Open-EQA[21] Habitat ScanNet/HM3D 2.53 0.27 Manual
MT-HM3D (Ours) Habitat HM3D 3.66 0.947 VLMs
B.2 Metrics
For multiple-choice questions, we have the model directly provide a predicted option, and then calculate the success rate (%). For open-vocabulary questions, we use two metrics to evaluate performance of the model: GPT score and ROUGEL [19]. The specific calculation methods are as follows:
@@ B.2 Metrics
where $LCS(\cdot,\cdot)$ is longest common subsequence, $C$ is prediction sentence, $R$ is ground-truth sentence, $cos$ is cosine similarity, ||$\cdot$|| ⋅ | is calculating the number of words.
@@
B.3 Prompts
We sample multiple images from the scene, stitch them together, and input the images along with a prompt 8 to generate proposed question-answer pairs.
1. do not ask "why" and "how";
2. the names of the objects must appear in the question;
3. avoid descriptions such as "left image" or "right image";
For example:
Does <OBJ1> share same color as <OBJ2> in <ROOM>?
Does <OBJ1> in <ROOM1> share same color as <OBJ2> in <ROOM2>?
Is <OBJ1> bigger/smaller than <OBJ2> in <ROOM>?
Is <OBJ1> in <ROOM1> bigger/smaller than <OBJ2> in <ROOM2>?
Is <OBJ1> closer than/farther from <OBJ2> than <OBJ3> in <ROOM>?
Is <ROOM1> bigger/smaller than <ROOM2> in the house?
How many <OBJ1> in <ROOM1>?
How many rooms have <OBJ1>?
The output must strictly following the format below:
Objects: [object1]; [object2]
Question: [question]
Options: [A. xxx; B. xxx; C. xxx; D. xxx]
Answer: [answer]
Label: [label]
Prompt 8: Prompt for data generation.
After obtaining the question-answer pair, we use prompts 9 to verify the validity of the question and answer.
(1) The question requires additional visual input to answer;
The question is as follows: {}
Prompt 9: Prompt for Verify the rationality of the data.
Appendix C More Experiments
@@
max retrieval num 10
@@
C.1 Implementation Details
We have listed all the hyper-parameters used for our best results in Table  5.
C.2 Parameter Analysis
As shown in Table  6, the selection of the K value is a pivotal factor, as both excessively large and unduly small K values can precipitate suboptimal performance. An insufficiently small K value engenders a dearth of memorized information, whereas an inordinately large K value not only induces informational redundancy but also fosters hallucinations in the response module. We believe that using different K values for different problems may be a more optimal strategy. Specifically, a dynamic memory window, rather than a fixed K value, could better adapt to varying complexities, which warrants further investigation.
@@ C.2 Parameter Analysis
C.3 Real-World Results
We conducted experiments using the Unitree Go2 robot in both office settings and artificially constructed scenarios. The RGB-D images were captured using the RealSense D435i depth camera, with DDS serving as the underlying middleware, and the control of the agent was accomplished on the ROS2 platform [20]. The multi-modal large model was deployed on 8 * Nvidia Tesla L40.
As shown in Figure  6, in the handmade scenario, we posed an open-vocabulary question to the agent: “How many apples are there in the scene?" After a thorough exploration, the agent provided an answer and did not recount when revisiting previously explored areas. In the office setting, we inquired, “Is the black garment placed on the chair?" The agent swiftly deduced the correct answer following a simple planning process.
@@ C.3 Real-World Results
C.4 Visualization
As shown in Figure  7, we have visualized ExploreEQA and MemoryEQA to demonstrate the advantages of our global and local memory mechanisms. The figure above illustrates an example of local memory. The baseline method provided an incorrect answer immediately after observing a sofa, whereas MemoryEQA, upon observing the first sofa, did not retrieve information about another sofa from its memory and thus continued its exploration, ultimately delivering the correct response.

### v2

@@ 5.3 Ablation Studies
As shown in Table 3, we evaluated different strategies’ impact on MemoryEQA performance. The k-value selection depends on memory retrieval: without retrieval, all memories are used as input, and without k-value selection, k is fixed at 5. Initially, the baseline model without memory strategies performs poorly, indicating suboptimal memory management. Introducing memory update (U) shows slight improvement, but its effect on success and efficiency is limited, likely due to low memory update needs in static scenarios. When memory retrieval (R) is combined with memory update (U+R), significant improvements are observed, enabling the system to access more relevant past experiences, thus boosting performance. The most significant gains occur when all three strategies are applied, with k-value selection enhancing retrieval by removing redundant information and selecting relevant data. Overall, combining memory update, retrieval, and k-value selection leads to the most efficient memory management.
As shown in Table 4, the ablation study evaluates the impact of the memory information into various module on MT-HM3D and HM-EQA datasets. Without memory (None), baseline success rates are 30.22% (MT-HM3D) and 51.5% (HM-EQA), with suboptimal step efficiency. Introducing the Stop Criterion (S) alone yields marginal improvements in MT-HM3D (+0.69% Succ., reduced steps). Combining Stopping and Answering Module (S+A) significantly boosts performance: MT-HM3D’s success rises by 10.77% (to 40.99%), and HM-EQA improves by 3.21% (to 54.76%), confirming that memory-enhanced answering critically resolves ambiguities in both tasks. Full integration (S+A+P) achieves peak results: MT-HM3D reaches 41.95% Succ. (11.73% gain over baseline) with a 0.4 norm step reduction, while HM-EQA attains 56.65% success rate (+5.1%) and fewer steps (0.49). The planner’s contribution is crucial, in synergy with S and A to optimize task decomposition and trajectory planning. This demonstrates that memory benefits all modules cumulatively, with the planner driving efficiency by mitigating redundant exploration.
5.4 Visualization
Figure 3 illustrates the critical role of memory in embodied question answering (EQA) systems. In EQA tasks, an agent interacts with its environment while simultaneously processing information to answer questions. Memory is essential for retaining key details, such as the location of objects or the layout of the environment, which directly informs the agent’s decisions and actions. In the example shown, the agent uses its memory to track the positioning of objects—such as the sofa in different rooms—allowing it to answer the question of whether the sofas in the living room and bedroom are the same color. By leveraging memory, the system can make informed, context-sensitive decisions, navigating the environment and adjusting its actions to retrieve and analyze relevant information. This capability is fundamental for improving the accuracy and efficiency of embodied question answering tasks in dynamic, real-world scenarios.
@@ 6 Conclusion
In this paper, we propose a memory-centric EQA framework, named MemoryEQA. MemoryEQA establishes a comprehensive memory management strategy that includes storage, updating and retrieval, enabling memory information to more efficiently and accurately assist other modules in reasoning. Additionally, we have constructed a benchmark MT-HM3D based on the real-world HM3D scenes. This benchmark includes four categories of complex tasks, allowing for comprehensive validation of the memory capabilities of EQA models. The effectiveness of MemoryEQA has been demonstrated through performance on the MT-HM3D, HM-EQA and Open-EQA datasets.
[…]
@@
7 Method Details
$q_{p}$, $q_{s}$, and $q_{a}$ denote the queries used to retrieve memory information required by the planner, stopping module, and answering module, respectively. The specific form of the query is as follows
Retrieve memory entries describing the spatial layout and previously observed regions near the current viewpoint.
Current viewpoint: <agent_pose>
Visible objects: <object_list>
Goal type: <goal_type>
Return memory that contains navigable regions, unexplored frontiers, and previously seen locations relevant to planning the next movement.
Prompt 1: The query $q_{p}$ used by planner.
Retrieve memory entries related to the existence, location, or appearance of the target object or target region.
Current observation: <observation_text>
Target specification: <question_text>
Return memory that may indicate the target has been found or is likely within the currently observed region.
Prompt 2: The query $q_{s}$ used by stopping module.
Retrieve memory entries that contain semantic information relevant to answering the question.
Question: <question_text>
Current observation: <observation_text>
Return memory that includes object attributes, spatial relations, and scene facts related to the queried concept.
Prompt 3: The query $q_{a}$ used by answering module.
Figure 4: Statistic of our proposed MT-HM3D.
We utilize prompt LABEL:lst:scene_cap and prompt LABEL:lst:obj_cap to prompt VLM to generate coarse-grained scene description and fine-grained object description respectively.
@@ 7 Method Details
Prompt 4: Prompt for coarse-grained scene description.
Dataset Simulator Source Real Scenes Objects Region Creation
EQA-v1[das2018embodied] House3D SUNCG - - Rule-Based
MP3D-EQA[wijmans2019embodied] MINOS MP3D - - Rule-Based
MT-EQA[yu2019multi] House3D SUNCG 3.20 1.0 Rule-Based
IQA[gordon2018iqa] AI2THOR - - - Rule-Based
EXPRESS-Bench[jiang2025beyond] Habitat HM3D 3.21 1.09 VLMs
HM-EQA[ren2024explore] Habitat HM3D 2.87 1.0 VLMs
Open-EQA[majumdar2024openeqa] Habitat ScanNet/HM3D 2.53 1.0 Manual
MT-HM3D (Ours) Habitat HM3D 3.44 1.49 VLMs
@@
Prompt 5: Prompt for fine-grained object description.
@@ E. Very high
Prompt 6: Prompt for the evaluation of confidence.
Prompt LABEL:lst:close_response and prompt LABEL:lst:open_response are used to prompt the answering module to closed-vocabulary response and open-vocabulary response.
{Question} Answer with the option’s single letter from the given choices directly.’
Prompt 7: Prompt for the response.
@@
Prompt 8: Prompt for the response on open-vocabulary data.
Next, prompt LABEL:lst:local_sem and prompt LABEL:lst:global_sem follow ExploreEQA [ren2024explore] setting to prompt VLM to output explore direction.
@@
Prompt 9: Prompt for the explore direction.
@@
Prompt 10: Prompt for the Exploring Value.
8 Data Construction
8.1 Statistics
To evaluate the memory capabilities of EQA models, we propose an EQA dataset that features multiple targets distributed across various regions. As shown in Table  5, our dataset exhibits the highest average number of targets and regions, significantly increasing the complexity of the task. The requirement to explore multiple targets across different regions poses a substantial challenge to the model’s memory capacity. The statistical details of the MT-HM3D dataset are illustrated in Figure  4.
8.2 Metrics
For multiple-choice questions, we have the model directly provide a predicted option, and then calculate the success rate (%). For open-vocabulary questions, we use two metrics to evaluate performance of the model: GPT score and ROUGEL [lin2004rouge]. The specific calculation methods are as follows:
@@ 8.2 Metrics
where $LCS(\cdot,\cdot)$ is longest common subsequence, $C$ is prediction sentence, $R$ is ground-truth sentence, coscos is cosine similarity, $|\cdot|$ is calculating the number of words.
@@
max retrieval num 10 $\beta$ 1.0
$\beta_{p}$ 5.0 $\beta_{r}$ 15.0
$\alpha$ 0.5 $\alpha_{e}$ 1.0
@@
8.3 Prompts
We sample multiple images from the scene, stitch them together, and input the images along with a prompt LABEL:lst:data_gen to generate proposed question-answer pairs.
1. do not ask "why" and "how";
2. the names of the objects must appear in the question;
3. avoid descriptions such as "left image" or "right image";
For example:
Does <OBJ1> share same color as <OBJ2> in <ROOM>?
Does <OBJ1> in <ROOM1> share same color as <OBJ2> in <ROOM2>?
Is <OBJ1> bigger/smaller than <OBJ2> in <ROOM>?
Is <OBJ1> in <ROOM1> bigger/smaller than <OBJ2> in <ROOM2>?
Is <OBJ1> closer than/farther from <OBJ2> than <OBJ3> in <ROOM>?
Is <ROOM1> bigger/smaller than <ROOM2> in the house?
How many <OBJ1> in <ROOM1>?
How many rooms have <OBJ1>?
The output must strictly following the format below:
Objects: [object1]; [object2]
Question: [question]
Options: [A. xxx; B. xxx; C. xxx; D. xxx]
Answer: [answer]
Label: [label]
Prompt 11: Prompt for data generation.
After obtaining the question-answer pair, we use prompts LABEL:lst:data_verify to verify the validity of the question and answer.
(1) The question requires additional visual input to answer;
The question is as follows: {}
Prompt 12: Prompt for Verify the rationality of the data.
@@ 1. Select the question label from the following list: [Comparison, Relationship, Counting, Attribute];
9 More Experiments
9.1 Implementation Details
We have listed all the hyper-parameters used for our best results in Table  6.
9.2 Parameter Analysis
As shown in Table  7, the selection of the K value is a pivotal factor, as both excessively large and unduly small K values can precipitate suboptimal performance. An insufficiently small K value engenders a dearth of memorized information, whereas an inordinately large K value not only induces informational redundancy but also fosters hallucinations in the response module. We believe that using different K values for different problems may be a more optimal strategy. Specifically, a dynamic memory window, rather than a fixed K value, could better adapt to varying complexities, which warrants further investigation.
9.3 Real-World Results
We conducted experiments using the Unitree Go2 robot in both office settings and artificially constructed scenarios. The RGB-D images were captured using the RealSense D435i depth camera, with DDS serving as the underlying middleware, and the control of the agent was accomplished on the ROS2 platform [macenski2022robot]. The multi-modal large model was deployed on 8 * Nvidia Tesla L40.
As shown in Figure  6, in the handmade scenario, we posed an open-vocabulary question to the agent: “How many apples are there in the scene?” After a thorough exploration, the agent provided an answer and did not recount when revisiting previously explored areas. In the office setting, we inquired, “Is the black garment placed on the chair?” The agent swiftly deduced the correct answer following a simple planning process.
9.4 Visualization
As shown in Figure  5, we have visualized ExploreEQA and MemoryEQA to demonstrate the advantages of our global and local memory mechanisms. The figure above illustrates an example of local memory. The baseline method provided an incorrect answer immediately after observing a sofa, whereas MemoryEQA, upon observing the first sofa, did not retrieve information about another sofa from its memory and thus continued its exploration, ultimately delivering the correct response.